Solve two dimensional Laplace equation d²u/dx²+d²u/dy²=0 ,subject to the condition u(0,y)=u(l,y)=u(x,0)=0, u(x,a)=sin nmx/l


1
Expert's answer
2021-07-05T17:56:25-0400

Given equation-


"\dfrac{d^{2}u}{dx^{2}}+\dfrac{d^{2}u}{dy^{2}}=0".........1)


, subject to condition u(0,y)=u(l,y)="u(x,0)=0,u(x,a)=sin\dfrac{nmx}{l}"



Let "u=XY"


where X is a function of x only and Y is a function of y only .


"=\dfrac{\partial^{2}u}{\partial x^{2}}=\dfrac{\partial(XY)}{\partial x^{2}}=Y\dfrac{d^{2}X}{dx^{2}}"


"=" "\dfrac{\partial^{2}u}{\partial y^{2}}=\dfrac{\partial(XY)}{\partial y^{2}}=X\dfrac{d^{2}Y}{dy^{2}}"


From 1) ,we can say that ,


"YX''+XY''=0"


"\implies" "\dfrac{X''}{X}+\dfrac{Y''}{Y}=0"


Case 1, we have


"\dfrac{X''}{X}=" "-\dfrac{Y''}{Y}=p^{2}"


"(i)" "\dfrac{X''}{X}=p^{2}"


"X''-P^{2}X=0"


auxiliary equation is given by -


"m^{2}-p^{2}=0"


"m={\pm}p"


"CF=C_1e^{px}+C_2e^{-px}"


"PI=0"


"X=C_1e^{px}+C_2e^{-px}"


"(ii)"

"\dfrac{-Y''}{Y}=p^{2}\implies" "Y''+p^{2}Y=0"



Auxiliary equation is "m^{2}+p^{2}=0\implies" "m={\pm}pi"


"CF=C_3cos py+C_4sinpy"


"PI=0"


"y=" "C_3cos py+C_4sinpy"


Now X(0)=0


"c_1+c_2=0\implies" "c_2=-c_1"


"X(l)=0"


"\implies" "C_1e^{pl}++C_2e^{-pl}=0"


"\implies" "C_1(e^{pl}-e^{-pl})=0"


"c_1=0"

"c_2=0"


X=0"\implies" U=XY =0 which is impossible.


hence we reject case 1 .


case 2) , when



"\dfrac{X''}{X} =-\dfrac{Y''}{Y}=0"


"(i)" "\dfrac{X''}{X}=0"


"X''=0\implies" "X=c_5x+c_6"


"(ii)" "-\dfrac{Y''}{Y}=0"


"\implies" "Y''=0" "\implies" "Y=c_7y+c_8"



Now , "X(0)=0\implies" "c_6=0"


"X(l)=0"


"\implies" "c_5l+c_6=0\implies" "c_5l=0"


"\implies" "c_5=0" (since "{\neq}0)"


"\therefore" "X=0"


Hence , we reject case 2.



Case 3)


"\dfrac{X''}{X}=-\dfrac{Y''}{Y}=p^{2}"


"(i)-\dfrac{X''}{X}=p^{2}"


"\implies" "X''+p^{2}X=0\implies" "\dfrac{d^{2}X}{dx^{2}}+p^{2}X=0"


Auxiliary equation is "m^{2}+p^{2}=0" =0 "m{\neq}pi"


"CF=c_9cos px+c_{10}sinpx"



"PI=0"


"X=" "c_9cos px+c_{10}sinpx"



"(ii)" "\dfrac{-Y''}{Y}=-P^{2}"



"\implies" "\dfrac{d^{2}Y}{dy^{2}}-p^{2}Y=0"



auxiliary equation is "m^{2}-p^{2}=0"



"\therefore" CF="C_{11}e^{PY}+C_{12}e^{-PY}"


"PI=0"


"\therefore" "CF=" "C_{11}e^{PY}+C_{12}e^{-PY}"


"PI=0"


"Hence ," "Y=C_{11}e^{py}+C_{12}e^{-py}"


Now, "X(0)=0\implies" "c_9=0"


"\therefore X=c_{10}sinpx"


"X(l)=0"


"c_{10}sinpl=0"


"\implies" "sinpl=0=sinn{\pi},n\isin" I


"\therefore" "p=\dfrac{n{\pi}}{l}"

"\therefore" "X=c_{10}sin\dfrac{n{\pi}x}{l}"


"Again ,Y(0)=0"


"\implies c_{11}+c_{12}=0\implies c_{11}=-c_{12}"


"Y=c_{11}(e^{py}-e^{-py})=c_{11}(e^\dfrac{n{\pi}y}{l}-e^\dfrac{-n{\pi}y}{l})"



"\therefore" "u=XY" ="c_{11}c_{10}sin\dfrac{n{\pi}x}{l}(e^\dfrac{n{\pi}y}{l}-e^\dfrac{-n{\pi}y}{l})"



"or\ u(x,y)=b_nsin\dfrac{n{\pi}x}{l}(e^\dfrac{n{\pi}y}{l}-e^\dfrac{-n{\pi}y}{l})"



"now, u(x,a)=sin\dfrac{n{\pi}x}{l}=b_nsin\dfrac{n{\pi}x}{l}" "(e^\dfrac{n{\pi}a}{l}-e^\dfrac{-n{\pi}a}{l})"



"\implies" "b_n=\dfrac{1}{(e^\dfrac{n{\pi}a}{l}-e^\dfrac{-n{\pi}a}{l})}" "=\dfrac{1}{2sinh \dfrac{(n{\pi}a)}{l}}"